OpenStudy (anonymous):

x^3-8 x^2+2x+4 %(x^2-4) @HighlyEducated87

OpenStudy (anonymous):

i need help please @HighlyEducated87

OpenStudy (highlyeducated87):

Just to be sure is there suppose to be a % sign in the problem

OpenStudy (anonymous):

x^3-8 x^2+2x+4 (x^2-4)

OpenStudy (anonymous):

x^3-8 / x^2+2x+4 (x^2-4) is this the euation sorry

OpenStudy (highlyeducated87):

Answer x^3-4x^2+2x-16

OpenStudy (anonymous):

x^3-8 / x^2+2x+4 (x^2-4)

OpenStudy (anonymous):

@HighlyEducated87 i need help

OpenStudy (highlyeducated87):

You want me to explain it

OpenStudy (anonymous):

yes

OpenStudy (highlyeducated87):

First you write your problem out Then you multiply 4 times(x^2-4) That will give you 4x^2-16 So now you rewrite the problem x^3-8x^2+2x+4x^2-16 Now combine all like terms and if a term cannot be combined it stays the same -8x^2+4x^2 which is -4x^2 becase -8+4= -4 Now you have nothing else to combine So that leaves you with x^3-4x^2+2x-16 x^3 cannot be combined, 2x cannot be combined and -16 cannot be combined So your answer is: x^3-4x^2+2x-16

OpenStudy (anonymous):

i think x^3-8/x^2+2x+4 divided by (x^2-4) = (x^3-8) / ((x^2+2x+4)(x^2-4)) = ((x-2)(x^2+2x+4)) / ((x^2+2x+4)(x^2-4)) ... factorise x^3-8 = (x-2) / (x^2-4) ... divide numerator and denominator by x^2+2x+4 = (x-2) / ((x+2)(x-2)) ... factorise x^2 - 4 = 1 / (x+2) ... divide numerator and denominator by x-2

OpenStudy (highlyeducated87):

Why didn't you say that?